An X-ray photon with a wavelength of 0.933 nm strikes a surface. The emitted electron has a kinetic energy of 994 eV .

What is the binding energy of the electron in kJ/mol? [Note that KE = 1/2mv^2 and 1 electron volt (eV) = 1.602×10^−19J.]
Answer #1

KE = 994 eV

= 994 * 1.602*10^-19 J

= 1.59*10^-16 J

Given:

lambda = 9.33*10^-10 m

use:

E = h*c/lambda

=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(9.33*10^-10 m)

= 2.131*10^-16 J

Use:

Eo = E - KE

= 2.131*10^-16 J - 1.59*10^-16 J

= 5.41*10^-17 J

This is bonding energy for 1 electron

For 1 mol of electron

Eo = 6.022*10^23 * 5.41*10^-17 J/mol

= 3.26*10^7 J/mol

= 3.26*10^4 KJ/mol

Answer: 3.26*10^4 KJ/mol